OpenStudy (anonymous):

23. Evaluate (10-2)-2 22. Perform the indicated operations: √8 + 4√32 + √6 21. Simplify (8m5)3 (2m3)3

OpenStudy (anonymous):

23.(10-2)-2=8-2=6 22.\[\sqrt{8}+4\sqrt{32}+\sqrt{6}=\sqrt{4\times2}+ 4\sqrt{16\times2}+\sqrt{6}\] =\[2\sqrt{2}+16\sqrt{2}+\sqrt{6}=18\sqrt{2}+\sqrt{6}\]